Interdependence and the Gains from Trade (Mobile-Enabied) Graded Assignment | Read C Due 01.23.19 at 1245 AM Attempts: Keep the Highest:/3 2. Comparative and absolute advantage Alex and Becky are farmers. Each one owns a 12-acre plot of land. The following table shows the amount of produce per year on a given acre. Each farmer chooses all acres to land and barley on the rest. alfalfa and barley each farmer car whether to devote all acres to producing alfalfa or barley or to produce alfalfta on some of Alfalfa Barley Alex 18 Becky 28 On the following graph, use the blue line (circle symbol) to plot Alexs production possibilities frontier (PPF), and use the purple line (diamond symbol to plot Beckys PPF 120 108
Alexs PPF Beckys PPp 0 60 120 180 240 300 360 420 480 540 00 ALFALFA (Bushels) has an absolute advantage in the production of alfalfa, andhas an absolute advantage in the production of barley Alexs opportunity cost of producing 1 bushel of barley is barley is bushels of alfalfa. Because Alex has aopportunity cost of producing comparative advantage in the production of barley, and has a bushels of alfalfa, whereas Beckys opportunity cost of producing 1 bushel of barley than Becky,has a comparative advantage in the production of alfalfa. NoticesTerms of Use Privacy NoticeSecurity Not
Answer #1

Alex's total production of Alfalfa = 18*12 = 216

Alex's total production of Barley = 6*12 = 72

Becky's total production of Alfalfa = 28*12 = 336

Becky's total production of Barley = 7*12 = 84

90 84 72 30 29 12 PPF AIFA LFA 332 Beekys has absoluts advantage in both bwey and alalla lexs opp cost q producing 1 bushel- 216 3 bshe 72 tost 30 89 Because Alex has a Jouwe oppoctunity Cost o pmdueing banky than Be Alex has a comparative ad ane Beuk anative advant, e-in-the has a pond uchm a.

Becky has an absolute advantage in the production of alfalfa and Becky has an absolute advantage in the production of barley. Alex's opportunity cost of producing 1 bushel of barley is 3 bushels of alfalfa, whereas Becky's opportunity cost of producing 1 bushel of barley is 4 bushels of alfalfa.

Because Alex has a lower opportunity cost of producing than Becky, Alex has a comparative advantage in the production of barley , and Becky has a comparative advantage in the production of alfalfa.